OpenStudy (fanduekisses):

If 4.0 g of oxygen are reacted with 2.0 g of hydrogen to produce water and leftover hydrogen, what mass of hydrogen is leftover? What is the mass of water produced? On what natural laws are your responses based?

OpenStudy (whpalmer4):

how many moles of oxygen and hydrogen are present?

OpenStudy (fanduekisses):

How do you tell how many moles?

OpenStudy (fanduekisses):

Oh atomic mass?

OpenStudy (whpalmer4):

Yes, you know that 1 mole of H atoms weighs 1.00794 grams, and 1 mole of O atoms weighs 15.9994 grams. Figure out which atom is in shorter supply (remembering that you need 2 H for every 1 O); that's the one which will determine how many moles of water you produce.

OpenStudy (rushwr):

\[2H _{2} +O _{2} \rightarrow 2H _{2}O\] As we can see 2 moles of hydrogen will react with one mole of oxygen . Since we are given with the masses of oxygen and hydrogen , 1st we will find the no. of moles of each present. We know no. of moles (n) = mass(m) divided by the molar mass (M) When u round off the molar mass of oxygen you can get it approximately around 16 OXYGEN --------> \[n = \frac{ 4g }{16gmol ^{-1}} = 0.25moles\] Hydrogen's molar mass is approximately 1 HYDROGEN-----------> \[n = \frac{ 2g }{ 1gmol ^{-1} } = 2mol\] According to the equation if we have 0.25 moles of oxygen then it will react with 0.5moles of hydrogen right (0.25*2) . Recall the reacting ratio. But no we can see that we have more hydrogen. To complete the reaction we only need 0.5 moles of hydrogen , so we have excess hydrogen . The left over moles of hydrogen = 2moles - 0.5 moles= 1.5 moles Use the same equation and subject the mass n= m/M m= n*M so mass of hydrogen left = 1.5 moles * 1gmol^-1 = 1.5g According to the reaction the stoichiometric no. of water formed is equal to that of H2 So if 0.5 moles of hydrogen is formed then 0.5 moles of water is formed. Now we know the moles of water (0.5moles) and the molar mass of water is around 18gmol^-1 Use the same equation m=n*M m = 18 * 0.5 = 9g

OpenStudy (rushwr):

Hope this will help you :)